# Fine-tuning Neural Network Quantum States

### v2

@@ Abstract
Recent progress in the design and optimization of Neural-Network Quantum States (NQS) have made them an effective method to investigate ground-state properties of quantum many-body systems. In contrast to the standard approach of training a separate NQS from scratch at every point of the phase diagram, we demonstrate that the optimization of a NQS at a highly expressive point of the phase diagram (i.e., close to a phase transition) yields interpretable features that can be reused to accurately describe a wide region across the transition. We demonstrate the feasibility of our approach on different systems in one and two dimensions by initially pretraining a NQS at a given point of the phase diagram, followed by fine-tuning only the output layer for all other points. Notably, the computational cost of the fine-tuning step is very low compared to the pretraining stage. We argue that the reduced cost of this paradigm has significant potential to advance the exploration of condensed matter systems using NQS, mirroring the success of fine-tuning in machine learning and natural language processing.
@@
Introduction. Over the last decade, neural networks have emerged as the most important general-purpose machine learning tool [1, 2]. Their versatility is evident in the most recent generation of neural networks based on the Transformer architecture [3], which was initially designed for natural language processing, and is now achieving state-of-the-art performance in fields as diverse as text generation [4, 5], computer vision [6], and protein contact prediction [7]. The success of deep neural networks is generally attributed to their ability to learn relevant features directly from data [1, 2]. This automatic feature discovery replaces the classic approach of machine learning, where one first designs a set of well-suited features to describe the inputs, and then trains a simple machine learning algorithm (e.g., linear regression) on top of those features to perform a given task. While hand-crafting suitable features is feasible for simple problems, the challenge arises in determining the optimal features for more complicated tasks. Feature learning methods based on neural networks instead learn good representations to solve the target task directly from raw data, outperforming hand-crafted representations in a variety of domains.
@@
In the last few years, neural networks are also increasingly used in condensed matter physics to approximate low-energy states of many-body quantum systems, both for spin and fermionic models [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 9]. In this context, Neural-Network Quantum States (NQS) parametrize the amplitude of a variational state |$\ket{\Psi_{\theta}}$ expandend in a proper basis {|$\{\ket{\bm{\sigma}}\}$, mapping input physical configurations $\bm{\sigma}$ to complex numbers $\braket{\bm{\sigma}}{\Psi_{\theta}}=\Psi(\bm{\sigma};{\theta})$. Within the Variational Monte Carlo framework the parameters ${\theta}$ of the state are optimized to minimize the variational energy $E_{\theta}=\braket{\Psi_{\theta}}{\hat{H}}{\Psi_{\theta}}/\braket{\Psi_{\theta}}{\Psi_{\theta}}$ [19].
@@
Figure 2: a) Ising chain. A ViT with hyperparameters $h=12$, $d=72$, $n_{l}=4$ is pretrained at $g/\Gamma=1$, on a chain with $N=100$ sites. After the fine-tuning, the square magnetization order parameter is computed and compared to DMRG results (bond dimension $\chi=10^{3}$). b) Heisenberg $J_{1}$-$J_{2}$ chain. A ViT with hyperparameters $h=12$, $d=192$, $n_{l}=4$ is pretrained at $J_{2}/J_{1}=0.4$, on a chain with $N=100$ sites. After the fine-tuning, the dimer order parameter is computed and compared to DMRG results ($\chi=10^{3}$). Inset : Relative error $\Delta\varepsilon$ (with respect to DMRG) of the same fully-connected network (RBM) trained on the hidden representations generated by the pretrained ViT and directly on configurations. c) Two dimensional Heisenberg $J_{1}$-$J_{2}$. A ViT with hyperparameters $h=18$, $d=216$, $n_{l}=8$ is pretrained at $J_{2}/J_{1}=0.5$, on a $6\times 6$ square lattice. After the fine-tuning, the structure factors at $\bm{k}=(\pi,\pi)$ and $\bm{k}=(0,\pi)$ are computed and compared to exact diagonalization results.
Viteritti et al. [20] recently introduced a new parametrization of NQS which explicitly leverages the feature learning perspective. In this viewpoint, a deep neural network is used to construct a map from the space of the physical configurations to abstract representations in a feature space, where the determination of the low-energy properties of the systems is simplified, and then a simple shallow network transform these representations into complex numbers. The resulting architecture achieved state-of-the-art performance on one of the most famous benchmark problems of the field [21].
In this Letter, we address a key conceptual question that this approach raises: given a system that exhibits a phase transition, do the representations learned to approximate the ground state near the transition point generalize to other points of the phase diagram? This question holds significance from both theoretical and practical perspectives. On the one hand, it has the potential to enhance the interpretability of Neural-Network Quantum States. On the other hand, it offers a practical advantage by circumventing the need to optimize the wave function from scratch for each point in the phase diagram.
Methods. Mirroring the feature learning perspective, we represent the neural quantum state as the composition of two functions [20]:
@@
where we have partitioned the variational parameters into two blocks $\theta=\{\phi,W\}$. The function $V(\bm{\sigma};\phi)$ is parameterized through a deep neural network, mapping physical configurations $\bm{\sigma}$ into a feature space, thereby generating for each input configuration $\bm{\sigma}$ a hidden representation $\bm{z}(\bm{\sigma})\in\mathbb{R}^{d}$, with $d$ the dimension of the feature space (an adjustable hyper-parameter of the network). Conversely, $f(\bm{z};W)$ is a shallow fully-connected neural network used to generate a single scalar value $f(\bm{z};W)\in\mathbb{C}$ from the hidden representations $\bm{z}$.
Given a system undergoing a phase transition, we want to investigate whether the representations learned near the transition point generalize to other points of the phase diagram. We perform the following experiment in two steps, as illustrated in Fig. 1:
Figure 3: Dimensional reduction of the hidden representations for a set of $M=3000$ configurations built using a ViT pretrained at $g/\Gamma=1$ with hyperparameters $h=12$, $d=72$, and $n_{l}=4$ for a cluster of $N=100$ sites. The data points represent UMAP projections of vectors $\bm{z}$. Both the colors and sizes of the points are related to the amplitudes predicted after the fine-tuning procedure at three distinct points along the phase diagram: ordered phase $g/\Gamma=0.4$ (left panel), transition point $g/\Gamma=1$ (central panel) and disordered phase $g/\Gamma=1.6$ (right panel).
@@
    Using the features constructed by the deep network (thus fixing its variational parameters ${\phi}_{p}$) we fine-tune the model by optimizing only the parameters $W$ of the output layer to approximate the ground states in the other points of the phase diagram, before and after the phase transition.
The choice of pretraining the model near the phase transition is related to the fact that we expect the ground state to be very entangled close to critical points, thus spanning a large portion of the Hilbert space [22]. On the contrary, in trivial phases, where only a few configurations have non-zero amplitudes, the ability of pretrained networks to generalise away from these phases will be limited. We apply this procedure on finite clusters and measure physical properties (e.g., order parameters) of various systems exhibiting, in the thermodynamic limit, phase transitions of different nature. In all the cases, the features extracted during the pretraining stage, close to transition points, lead to excellent results after the fine-tuning in all the other points of the phase diagram [23]. This is an important conceptual result, that reflects how neural networks can capture the essential quantum fluctuations in the vicinity of a phase transition.
The outlined methodology is generally applicable to any deep neural network, but to be concrete, in the following we parameterize the function $V(\bm{\sigma};\phi)$ using a Vision Transformer (ViT) with real-valued parameters [20, 21]. This choice is suggested by the extreme flexibility of the Transformer architecture [3, 6], already used to achieve highly accurate results in various types of systems, both one- and two-dimensional [24, 25, 21, 20, 26, 27, 28]. The hyperparameters of this architecture are the number of heads $h$ of the Multi-Head Factored Attention Mechanism [29], the embedding dimension $d$ and the number of layers $n_{l}$ (a detailed description of the architecture is reported in Ref. [20]). In addition, the function $f(\bm{z},W)$ is defined as
@@
where the number of neurons $K$ is chosen to be equal to $d$ and $2\times d$ in the pretraining and in the fine-tuning steps, respectively. In order to describe non-positive ground states, the parameters $W=\{b_{\alpha},\bm{w}_{\alpha}\}_{\alpha=1}^{K}$ of the linear transformation in Eq. (2) are taken complex-valued. Here, we set $g(\cdot)=\log\cosh(\cdot)$, thus $f(\bm{z},W)$ represents the well-known Restricted Boltzmann Machine (RBM) introduced by Carleo and Troyer [8]. The crucial difference is that in this case it is not applied directly on physical configurations $\bm{\sigma}$, but instead on hidden representations $\bm{z}$ generated by the Transformer [20]. We remark that this framework offers a huge computational advantage, since it requires the costly optimization of the full architecture, including the feature extractor $V(\bm{\sigma};\phi)$, only once in the pretraining step. Then, with the addition of a minimal cost, the targeted optimization of the RBM can be used to obtain an accurate description of the physical properties of the system in a wide region across the transition point. In what follows, we focus on spin $S=1/2$ models on a lattice on clusters where numerically exact solutions are available for comparison. In this case $\Psi(\bm{\sigma};{\theta})$ refers to the amplitude of the variational state |$\ket{\Psi_{\theta}}$ in the computational basis having a definite spin value in the $z$ direction, i.e., {|$\{\ket{\bm{\sigma}}=\ket{\sigma_{1}^{z},\cdots,\sigma_{N}^{z}}\}$ with $\sigma_{i}^{z}=\pm 1$.
Results. We start by considering the one-dimensional Ising model in transverse magnetic field, described by the following Hamiltonian
@@
First, we pretrain the full architecture at the critical point $g/\Gamma=1$. Then, we fine-tune only the output layer at different values of the external field, from $g/\Gamma=0.4$ to $g/\Gamma=1.6$, i.e., in both ferromagnetic and paramagnetic phases. The results for $N=100$ with periodic-boundary conditions are shown in Fig. 2a), in comparison with density-matrix renormalization group (DMRG) [30] calculations (on the same cluster). The high level of accuracy clearly demonstrates that the fine-tuned network is extremely effective in the prediction of the order parameter. Remarkably, the fine-tuning procedure involves optimizing merely $6.6\%$ of the total parameters, which is ten times faster than optimizing the entire network and demands significantly less GPU memory (see Supplemental Material .1 for a detailed description of the GPU memory requirements). The remarkable fact is that, by exclusively adjusting the parameters of the output (fully-connected) layer and keeping the clusters of the hidden representation fixed, it is possible to effectively describe both ordered and disordered phases.
Figure 4: Graphical representation of the hidden representations for the $J_{1}$-$J_{2}$ Heisenberg chain. The data points, corresponding to a sample of $M=3000$ physical configurations, represent UMAP projections of vectors $\bm{z}_{p}$ generated by a ViT with hyperparameters $h=12$, $d=192$, and $n_{l}=4$, pretrained at the point $J_{2}/J_{1}=0.4$ for a system size of $N=100$. The depicted colors correspond to the predicted phases (00 or $\pi$) after fine-tuning at two specific points within the phase diagram: $J_{2}/J_{1}=0.1$ (left panel) and $J_{2}/J_{1}=0.6$ (right panel). The left panel reveals a close resemblance between the cluster structure identified during the pretraining at $J_{2}/J_{1}=0.4$ which match the Marshall sign rule.
@@
In the following, we want to gain insights into the learning process of the fine-tuning stage. For that, we sample a set of $M$ configurations $\{\bm{\sigma}_{i}\}\sim|\Psi(\bm{\sigma};\theta_{p})|^{2}$ from the pretrained network and show the corresponding amplitudes after the finetuning procedure (visualizing them on top of UMAP [31] projections of the hidden representations $\bm{z}_{p}(\bm{\sigma}_{i})$, for $i=1,\dots,M$), see Fig. 3. To highlight the differences, both color and size of each point are proportional to their amplitudes. At the transition point ($g/\Gamma=1$), the configurations with all parallel spins (either up or down along $z$) have the largest amplitude; other configurations, with a few spin flips have also considerable weights (see middle panel). In the ordered phase ($g/\Gamma=0.4$), only one of these fully-polarized configurations is “selected”, i.e., frequently visited along the Monte Carlo sampling, and the amplitudes for all other configurations are practically negligible (left panel). This effect is related to the difficulty of simple sampling (that performs local spin flips) to overcome the (large) barrier that separates the two ground states, which are almost degenerate on finite systems. By contrast, in the disordered phase ($g/\Gamma=1.6$), many configurations have similar amplitudes: the two fully-polarized configurations showing a reduced weight compared to all the others (right panel).

### v3

@@ Abstract
Recent progress in the design and optimization of neural-network quantum states (NQSs) has made them an effective method to investigate ground-state properties of quantum many-body systems. In contrast to the standard approach of training a separate NQS from scratch at every point of the phase diagram, we demonstrate that the optimization of a NQS at a highly expressive point of the phase diagram (i.e., close to a phase transition) yieldsfeatures that can be reused to accurately describe a wide region across the transition. We demonstrate the feasibility of our approach on different systems in one and two dimensions by initially pretraining a NQS at a given point of the phase diagram, followed by fine-tuning only the output layer for all other points. Notably, the computational cost of the fine-tuning step is very low compared to the pretraining stage. We argue that the reduced cost of this paradigm has significant potential to advance the exploration of strongly-correlated systems using NQS, mirroring the success of fine-tuning in machine learning and natural language processing.
I Introduction
@@
Over the last decade, neural networks have emerged as the most important general-purpose machine learning tool [1, 2]. Their versatility is evident in the most recent generation of neural networks based on the Transformer architecture [3], which was initially designed for natural language processing, and is now achieving state-of-the-art performance in fields as diverse as text generation [4, 5], computer vision [6], and protein contact prediction [7]. The success of deep neural networks is generally attributed to their ability to learn relevant features directly from data [1, 2]. This approach replaces the classical one, where one first designs a set of well-suited features to describe the inputs, and then trains a simple machine learning algorithm (e.g., linear regression) to perform a given task. However, hand-crafting suitable features is feasible only for simple problems and becomes unfeasible when dealing with complicated tasks. By contrast, deep-learning methods learn good representations to solve the target task directly from raw data, outperforming hand-crafted representations in a variety of domains.
@@
In the last few years, neural networks are also increasingly used in condensed matter physics to approximate low-energy states of many-body quantum systems, for both spin and fermionic models [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 9]. In this context, neural-network quantum states (NQSs) parametrize the amplitude of a variational state |$\ket{\Psi_{\theta}}$ expandend in a proper basis {|$\{\ket{\bm{\sigma}}\}$, mapping input physical configurations $\bm{\sigma}$ to complex numbers $\braket{\bm{\sigma}}{\Psi_{\theta}}=\Psi(\bm{\sigma};{\theta})$. Within the variational Monte Carlo framework, the parameters ${\theta}$ of the state are optimized to minimize the variational energy $E_{\theta}=\braket{\Psi_{\theta}}{\hat{H}}{\Psi_{\theta}}/\braket{\Psi_{\theta}}{\Psi_{\theta}}$ [19].
Recently, a new parametrization of NQS which explicitly leverages the feature learning perspective has been introduced in Ref. [20]. In this viewpoint, a deep neural network is used to construct a map from the space of the physical configurations to abstract representations in a feature space, where the determination of the low-energy properties of the systems is simplified, and then a simple shallow network transform these representations into complex numbers. The resulting architecture achieved state-of-the-art performance on one of the most famous benchmark problems of the field [21].
In this work, we address a key conceptual question that this approach raises: given a system that exhibits a phase transition, do the representations that have learned to approximate the ground state near the transition point generalize to other points of the phase diagram? This question holds significance not only from a theoretical perspective but also from a practical one, as it provides a concrete advantage of avoiding the need of optimizing the wave function from scratch for each point in the phase diagram.
II Methods
@@
Mirroring the feature learning perspective, we represent the NQS as the composition of two functions [20]:
@@
where we have partitioned the variational parameters into two blocks $\theta=\{\phi,W\}$. The function $V(\bm{\sigma};\phi)$ is parameterized through a deep neural network, mapping physical configurations $\bm{\sigma}$ into a feature space, thereby generating for each input configuration $\bm{\sigma}$ a hidden representation $\bm{z}(\bm{\sigma})\in\mathbb{R}^{d}$, with $d$ the dimension of the feature space (an adjustable hyperparameter of the network). Conversely, $f(\bm{z};W)$ is a shallow fully connected neural network used to generate a single scalar value $f(\bm{z};W)\in\mathbb{C}$ from hidden representations $\bm{z}$.
Given a system undergoing a phase transition, we want to investigate whether the representations learned near the transition point may be generalized to other points of the phase diagram. We perform the following experiment in two steps, as illustrated in Fig. 1:
Figure 2: (a) Ising chain. A ViT with hyperparameters $h=12$, $d=72$, $n_{l}=4$ is pretrained at $g/\Gamma=1$, on a chain with $N=100$ sites. After the fine-tuning, the square magnetization order parameter is computed and compared to DMRG results (bond dimension $\chi=10^{3}$). (b) Heisenberg $J_{1}$-$J_{2}$ chain. A ViT with hyperparameters $h=12$, $d=192$, $n_{l}=4$ is pretrained at $J_{2}/J_{1}=0.4$, on a chain with $N=100$ sites. After the fine-tuning, the dimer order parameter is computed and compared to DMRG results ($\chi=10^{3}$). Inset : Relative error $\Delta\varepsilon$ (with respect to DMRG) of the same fully connected network (RBM) trained on the hidden representations generated by the pretrained ViT and directly on configurations. (c) Two dimensional Heisenberg $J_{1}$-$J_{2}$. A ViT with hyperparameters $h=18$, $d=216$, $n_{l}=8$ is pretrained at $J_{2}/J_{1}=0.5$, on a $6\times 6$ square lattice. After the fine-tuning, the structure factors at $\bm{k}=(\pi,\pi)$ and $\bm{k}=(0,\pi)$ are computed and compared to exact diagonalization results.
@@
    Using the features constructed by the deep network (thus fixing its variational parameters ${\phi}_{p}$) we fine-tune the model by optimizing only the parameters $W$ of the output layer to approximate the ground states in the other points of the phase diagram, on both sides of the phase transition.
The pretraining of the architecture is carried out near the critical point, where long-range correlations are present and may be established in the body of the NQS. Then, the last (shallow) output layer, which is fine-tuned in a different point in the phase diagram, can either reinforce correlations and establish true long-range order or weaken them and yield to a short-range state (or even keep the state critical). On the other hand, in trivial phases, where only a few configurations have non-zero amplitudes, the ability of pretrained networks to generalize away from these phases is limited (see Appendix A).
We apply this procedure on finite systems and measure physical properties (e.g., order parameters) of various systems exhibiting, in the thermodynamic limit, phase transitions of different nature. In all cases, the features extracted during the pretraining stage, close to transition points, lead to excellent results after fine-tuning at the other points of the phase diagram [22]. This result reflects how neural networks can capture the essential quantum fluctuations in the vicinity of a phase transition. We stress that this approach differs from the standard transfer-learning paradigm in which a neural network is initially trained to solve a specific task, and then all of the parameters are trained to solve a different task. To the best of our knowledge, fine-tuning experiments on NQS have not been explored previously.
The methodology outlined here is generally applicable to any deep neural network, but to be concrete, in the following we parametrize the function $V(\bm{\sigma};\phi)$ using a Vision Transformer (ViT) with real-valued parameters [20, 21]. This choice is suggested by the flexibility of the Transformer architecture [3, 6], already used to achieve highly accurate results in various types of systems, both one- and two-dimensional [23, 24, 21, 20, 25, 26, 27]. The hyperparameters of this architecture are the number of heads $h$ of the Multi-Head Factored Attention Mechanism [28], the embedding dimension $d$ and the number of layers $n_{l}$ (a detailed description of the architecture is reported in Ref. [20]). In addition, the function $f(\bm{z},W)$ is defined as
@@
where the number of neurons $K$ is chosen to be equal to $d$ and $2\times d$ in the pretraining and in the fine-tuning steps, respectively. In order to describe nonpositive ground states, the parameters $W=\{b_{\alpha},\bm{w}_{\alpha}\}_{\alpha=1}^{K}$ of the linear transformation in Eq. (2) are taken to be complex valued. Here, we set $g(\cdot)=\log\cosh(\cdot)$, thus $f(\bm{z},W)$ represents the well-known Restricted Boltzmann Machine (RBM) introduced by Carleo and Troyer [8]. The crucial difference is that, in this case, it is not applied directly on physical configurations $\bm{\sigma}$, but instead on hidden representations $\bm{z}$ generated by the Transformer [20]. We remark that this framework offers a huge computational advantage, since it requires the costly optimization of the full architecture, including the feature extractor $V(\bm{\sigma};\phi)$, only once in the pretraining step. Then, with the addition of a minimal cost, the targeted optimization of the RBM can be used to obtain an accurate description of the physical properties of the system in a wide region across the transition point. In what follows, we focus on spin $S=1/2$ models on a lattice considering system sizes where numerically exact solutions are available for comparison. In this case $\Psi(\bm{\sigma};{\theta})$ refers to the amplitude of the variational state |$\ket{\Psi_{\theta}}$ in the computational basis having a definite spin value in the $z$ direction, i.e., {|$\{\ket{\bm{\sigma}}=\ket{\sigma_{1}^{z},\cdots,\sigma_{N}^{z}}\}$ with $\sigma_{i}^{z}=\pm 1$.
Finally, we mention a few details on the Monte Carlo procedure adopted here. During the pretraining stage, we perform $N_{\text{opt}}=10000$ optimization steps. Then, during the fine-tuning stage, the number of steps is reduced to $N_{\text{opt}}=3000$. For all the simulations, we estimate stochastically the observables choosing a number of samples of $M=3000$. The optimization of the variational parameters is performed with the Stochastic Reconfiguration (SR) method [29]. In particular, working with variational states featuring approximately $P=10^{6}$ parameters, we employ the alternative formulation of SR [21, 30] efficient in the regime $P\gg M$ (available in NetKet [31] under the name of VMC_SRt.). We use a cosine decay scheduler for the learning rate, setting the initial value to $\tau=0.03$.
III Results
III.1 The one-dimensional quantum Ising model
We start by considering the one-dimensional Ising model in transverse magnetic field, described by the following Hamiltonian
@@ III.1 The one-dimensional quantum Ising model
First, we pretrain the full architecture at the critical point $g/\Gamma=1$. Then, we fine-tune only the output layer at different values of the external field, from $g/\Gamma=0.4$ to $g/\Gamma=1.6$, i.e., in both ferromagnetic and paramagnetic phases. The results for $N=100$ with periodic-boundary conditions are shown in Fig. 2(a), compared with density-matrix renormalization group (DMRG) [32] calculations (on the same system). The high level of accuracy demonstrates that the fine-tuned network is effective in the prediction of the order parameter. Remarkably, the fine-tuning procedure involves optimizing merely $6.6\%$ of the total parameters, which is ten times faster than optimizing the entire network and demands significantly less GPU memory (see Appendix B for a detailed description of the GPU memory requirements). The remarkable fact is that, by exclusively adjusting the parameters of the output (fully connected) layer and keeping the clusters of the hidden representation fixed, it is possible to effectively describe both ordered and disordered phases.
Figure 3: Dimensional reduction of the hidden representations for a set of $M=3000$ configurations built using a ViT pretrained at $g/\Gamma=1$ with hyperparameters $h=12$, $d=72$, and $n_{l}=4$ for a system of $N=100$ sites. The data points represent UMAP projections of vectors $\bm{z}$. Both the colors and sizes of the points are related to the amplitudes predicted after the fine-tuning procedure at three distinct points along the phase diagram: ordered phase $g/\Gamma=0.4$ (left panel), transition point $g/\Gamma=1$ (central panel) and disordered phase $g/\Gamma=1.6$ (right panel).
@@
In the following, we want to gain insights into the learning process of the fine-tuning stage. For that, we sample a set of $M$ configurations $\{\bm{\sigma}_{i}\}\sim|\Psi(\bm{\sigma};\theta_{p})|^{2}$ from the pretrained network and show the corresponding amplitudes after the finetuning procedure [visualizing them on top of UMAP  [^1] UMAP [56] is a general purpose dimension reduction technique for machine learning. It is constructed from a theoretical framework based in Riemannian geometry and algebraic topology, see Ref. [56] for more details. projections of the hidden representations $\bm{z}_{p}(\bm{\sigma}_{i})$, for $i=1,\dots,M$]; see Fig. 3. To highlight the differences, both color and size of each point are proportional to their amplitudes. At the transition point ($g/\Gamma=1$), the configurations with all parallel spins (either up or down along $z$) have the largest amplitude; other configurations, with a few spin flips have also considerable weights (see middle panel). In the ordered phase ($g/\Gamma=0.4$), only one of these fully polarized configurations is “selected”, i.e., frequently visited along the Monte Carlo sampling, and the amplitudes for all other configurations are practically negligible (left panel). This effect is related to the difficulty of simple sampling approaches (that performs local spin flips) to overcome the (large) barrier that separates the two ground states, which are almost degenerate on finite systems. By contrast, in the disordered phase ($g/\Gamma=1.6$), many configurations have similar amplitudes: the two fully polarized configurations showing a reduced weight compared to all the others (right panel). A brief discussion about the connection between the features extracted by the NQS and the order parameter is given in Appendix C.
III.2 The one-dimensional $J_{1}$-$J_{2}$ Heisenberg model